OpenStudy (anonymous):

HELP WITH ALGEBRA?!! WILL MEDAL + FAN!!! The length of a rectangle is 7 mm longer than its width. Its perimeter is more than 62 mm. Let w equal the width of the rectangle. The length has to be greater than ____ mm.

OpenStudy (anonymous):

you should already know that the perimeter is the length times 2 plus the width times 2. using this logic, you can construct a formula by substituting the information that they have given you. to make it easier to answer change the width to x 42=2x+2(x+5) expand the brackets- 42=2x+2x+10 collect the like terms- 4x-32=0 rearrange- 4x=32 remember your x is you width so the formula for the width would be width=8 to find the length all you need to do is add 5 to the width so the length is 13 if you want to check this to make sure it is correct just substitute the values into the formula for the perimeter= 2(lenght+width) 2(13+8) 26+16=42

OpenStudy (anonymous):

Thanks @champsionsaregreat!!!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

thx for the faning